OpenStudy (anonymous):

Solve for c. 2abc + d = 3 3 - d/2ab 3-d/2ab 3/2ab+d

OpenStudy (anonymous):

this would be the first one because u switch the sign and c for 3 right?

OpenStudy (anonymous):

do ab or d have values?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

first of all u have to bring whatever u do not want to isolate to the other side, what would b the step :P?

OpenStudy (anonymous):

u want to isolate c so ur first step is to bring d to the other side 3-d=2abc

OpenStudy (anonymous):

now think of 2ab as a constant, call it z so u have 3-d=zc

OpenStudy (anonymous):

how can u solve for c now?

OpenStudy (anonymous):

it becomes (3-d)/z=c =(3-d)/ab=c get it :)?

OpenStudy (anonymous):

yes;) that helped thanks

OpenStudy (anonymous):

no problem :)